40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 2604     Entry time: Thu Jul 15 23:37:53 2021
Author: Koji 
Type: Summary 
Category: Cryo vacuum chamber 
Subject: Bonding work for the prep of the preliminary suspension test 

[Stephen / Koji]

Bonding work for the prep of the preliminary suspension test

- 1" sq mirror-ish polished SUS piece was bonded to a face of the silicon mass. We chose the location right next to a line on the barrel. (Attachment 1)

- The mass was flipped with two more same thickness pieces used for the spacers to keep the mass horizontal.

- A pair of an OSEM and dumbbell-magnet was brought from the 40m (courtesy by Yehonathan). The magnet was glued on the mass at the opposite position of the attached mirror because the optical ports are going to be arranged to share an axis. A piece of cryo varnish was also painted with a piece of cigarette paper at the center of the mass so that we can attach an RTD. (Attachment 2)

Next Things To Do (Attachment 3)

  • Vent the chamber
  • We will move an optical port to the opposite position of the other port.
  • A DB9 feedthru is going to be installed.
     
  • Suspension
    • Move the sus frame in the chamber
    • Suspend the mass
  • Sensor arrangement
    • Set up the oplev
    • Hold the OSEM at the height of the magnet
    • Set up a camera to observe the magnet-OSEM clearance
    • We improvise the DB crimping sockets so that we can electrically connect the OSEM (optional)
  • Pump down / cool down the chamber
    • The main target of the cooling is to check the cooling capability of the test mass mainly with radiative cooling.
    • An optional target is to observe the misalignment as a function of the temperature -
      • -> Oplev signals are to be connected to CDS / check if CDS is logging the data
    • Check if the OSEM/magnets survive the thermal cycle
    • If possible we can try to actuate the OSEM / check the LED/PD function at the cryo temp
       
Attachment 1: P_20210715_170102-1.jpg  139 kB  Uploaded Fri Jul 16 00:38:48 2021  | Hide | Hide all
P_20210715_170102-1.jpg
Attachment 2: P_20210715_172218-1.jpg  77 kB  Uploaded Fri Jul 16 00:38:56 2021  | Hide | Hide all
P_20210715_172218-1.jpg
Attachment 3: experiment_plan.pdf  99 kB  Uploaded Fri Jul 16 00:55:19 2021  | Hide | Hide all
experiment_plan.pdf
ELOG V3.1.3-